1. Based on the data in the table below (a) what is the Kd for hormone binding by Protein1 and Protein ? (b) Which of these proteins binds most tightly to this hormone? Work this problem on a separate piece of graph paper. Y for Protein 2 ormone concentration (nM) Yfor Protein 1 0.17 0.33 0.5 0.8 0.91 0.95 0.98 0.2 .5 5T 4 10 20 50 0.5 0.67 0.89 0.95 0.97 0.99 2. Calculate the fractional saturation of hemoglobin at po, 20 torr and n -3. Let Pso 26 torr. 3. Calculate the Pso value for hemoglobin if Yo2 - 0.70 when po, 20 torr 02
Answer #1

Question 1: a) the Kd for any given protein binding with a ligand (in this case the hormone) is that concentration at which half the ligand binding sites are bound by the hormone, i.e., Y = 0.5. So, in the case of protein 1, the Kd is 1nM while in the case of protein 2, the Kd is 0.5nM.

b) As it is clear from the above table, the Kd for protein 2 is lesser than the protein 1, so protein 2 has more affinity towards the mentioned hormone (as a lesser concentration of hormone is required for filling the 50% ligand binding sites for protein 2). And thus we can say that protein 2 binds more tightly with the given hormone.

Question 2: pO2: 20 torr

n=3 and P50= 26 torr

You have written it right in the paper, we need to use Hill's equation to solve this problem.

Yo2=pO2n/(P50n+pO2n)

=203/(263+203)

=8000/(17576+8000)

=0.3127

Question 3: can be solved the similar way as question 2.

Yo2=pO2n/(P50n+pO2n)

0.70=203/(x3+203)

x3=(8000/0.70)-8000

x3=11428.57-8000

x3=3428.57

x or P50= 15.0789 torr